OpenStudy (anonymous):

need help with this question 4^3sqrt135

OpenStudy (thomas5267):

Is this the question you're asking? \[ 4^3\sqrt{135} \]

OpenStudy (anonymous):

yes i need help i dont understand how to do it

OpenStudy (thomas5267):

Calculate \(4^3\), then factorize 135.

OpenStudy (anonymous):

ok so 4^3 equals 64

OpenStudy (anonymous):

and 135 factorized is 3^3 x 5

OpenStudy (thomas5267):

Do you know that radials (square roots) have the following property? \[ \sqrt{3^3\times5}=3\sqrt{3\times5} \]

OpenStudy (anonymous):

no i did not know that, this confuses me so much

OpenStudy (thomas5267):

OK, radicals have the following properties. \[ \sqrt{xy}=\sqrt{x}\times\sqrt{y} \] Similarly, you can have something like this. \[ \sqrt{x^2y}=\sqrt{x^2}\times\sqrt{y}=x\times\sqrt{y}=x\sqrt{y} \]